OpenStudy (anonymous):

1. A 12-foot ladder is leaning against a wall that makes a 90-degree angle with the ground. A troublemaker is slowly pulling the bottom of the ladder along the ground, away from the wall at a rate of 2 feet per minute when the top of the ladder is 6 feet above the ground. b. At that moment, what is the rate of change of the angle between the ladder and the ground?

OpenStudy (anonymous):

b) Is similar to part a but instead of using Pythagorean equations use cos(θ)=b/12 h = 6\[b = \sqrt{144-6^2}= 6\sqrt{3}\] and \[\theta = \frac{\pi}{6}\] \[\frac{d}{dt}[cos(\theta )] = \frac{d}{dt}[\frac{b}{12}]\] \[-sin(\theta ) \frac{d\theta}{dt}= \frac{1}{12} \frac{db}{dt}\] since db/dt = 2. \[\frac{d\theta }{dt} = \frac{-1}{2sin(\theta)}2\] \[\frac{d\theta }{dt} = \frac{-1}{sin(\theta)}\] Hense , at h = 6. \[\frac{d\theta }{dt} = -2\]

OpenStudy (anonymous):

This is in radians BTW.^^^^^